OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 1) and (4, 13)? y – 1 = –2(x – 4) y – 1 = 2(x + 2) y – 13 = –2(x + 2) y – 13 = 2(x + 4)

OpenStudy (texaschic101):

do you know how to find the slope ?

OpenStudy (texaschic101):

@pamelasaucedo ...you there ?

OpenStudy (anonymous):

it's b

OpenStudy (lopezking1):

y=mx+b

OpenStudy (texaschic101):

let me check...hold on

OpenStudy (lopezking1):

Point-slope form looks like this: y - y1 = m (x - x1) The point (x1, y1) satisfies this equation because when (x,y) = (x1, y1) it becomes 0 = 0, which is always true. This equation describes a line of slope m because when x increases by 1, y increases by m. You can use either point, so both of these answers are correct: y – 1 = m(x + 2) y – 13 = m(x - 4) Without finding the slope you can already eliminate all choices except (b) because they don't fit either of these patterns. Calculate the slope between the two points by dividing the change in y by the change in x: m = (y2 - y1) / (x2 - x1) You can take the two points in either order, getting 12/6 or -12/-6. Either way, m = 2, confirming that the correct answer is (b).

OpenStudy (texaschic101):

m = (y2 - y1) / (x2 - x1) m = (13 - 1) / (4 - (-2) m = 12/6 = 2 y - y1 = m(x - x1) y - 1 = 2(x - (-2) y - 1 = 2x + 2 you are correct